OpenStudy (anonymous):

True or false. Linear Algebra. Each eigenvector of an invertible marix A is also an eigenvector of A^-1

jimthompson5910 (jim_thompson5910):

If k is an eigenvalue for the matrix A, then Ax = kx A^(-1)*Ax = A^(-1)*(kx) I*x = A^(-1)*(kx) x = A^(-1)*(kx) x = (A^(-1)x)*k (1/k)*x = A^(-1)x A^(-1)x = (1/k)*x So 1/k is the eigenvalue for the matrix A^(-1) So the statement "Each eigenvector of an invertible marix A is also an eigenvector of A^-1" is false (since k does not equal 1/k) Notice how this is only true if k = 1 or k = -1